# Interactive Inference: A Neuromorphic Theory of Human-Computer Interaction

### v3

@@ 3.2. Ten Postulates on Capacity and Error in Tasks
    Difficulty. The difficulty of the task equals the number of bits of relative entropy remaining to be processed. An uncompleted task has the full number of bits remaining. A fully accomplished task has a minimum bits of remaining entropy to be processed (preferably zero). Converting bits to time estimates requires a multiplication of this difficulty with an empirically derived distribution of time performance. E.g., the prior indicates prior statistical knowledge, built up over time, about where socks should be in a second task: that of obtaining a new pair of socks. After the task of organizing socks is complete, when the user wants to obtain any number of socks owned, the probability of obtaining them in the prior location is 100% because the posterior distribution of socks corresponds 100% to the prior, the location inside the drawer. By reducing the entropy in the distribution of socks in the first task to fit the goal, the user has been able to make the currently estimated location of socks, the posterior or progress distribution P, correspond to the memorized prior, the goal distribution G.
@@
    Goal and Progress Distributions. Goal and progress of a task are defined as probability distributions on the scale of task outcomes $s$. The goal distribution $G$ describes the prior, the probability of positive task outcomes on this scale. The progress distribution $P$ describes the posterior, the probability of being in a current position along the scale of task outcomes $s$. Each distribution has a mean value $\mu$ on this scale. For the goal distribution, the mean $\mu_{g}$ signifies the value with the highest probability of a correct task outcome. For the progress distribution, the mean $\mu_{p}$ denotes the value with the highest probability of the current task outcome. Each distribution also has an associated variance. The variance $\sigma_{g}^{2}$ or standard deviation $\sigma_{g}$ of the goal distribution $G$ describes the tolerance for error that is allowed in achieving goal $G$. The reciprocal of this variance $\frac{1}{\sigma_{g}^{2}}$ is the precision with which the task is to be completed. The variance $\sigma_{p}^{2}$ or standard deviation $\sigma_{p}$ of the progress distribution $P$ describes the speed with which the task outcome is moving towards the goal distribution: the speed of progress on the task. The reciprocal of this variance $\frac{1}{\sigma_{p}^{2}}$ is the precision with which the task is being executed.
    From postulate 5 it is becomes apparent that, in order for the goal and progress distributions to match, the user must not just reduce the distance between the means of progress and goal distributions to zero, but also match the precision (or variance) of both distributions. That is, a more precise goal will require, at some point during the task, a more precise movement on the scale of possible task outcomes towards that goal. In other words, to avoid error, the user will, at some point, need to move more slowly on the scale of possible task outcomes. This is typically towards the end of the task, when the user is closer to the goal. It also apparent that a not so precise goal definition allows faster progress, i.e., users can be more sloppy in the execution of the task. As such, the ratio between the difference of means and the standard deviation reflects a precision/accuracy trade-off (mistakenly known as speed/accuracy trade-off) that is inherent to all tasks. Because the two distributions will need to match on average, the individual variances of goal and progress distributions can be simplified as equal. This brings us to the next postulate:
@@
    Signal and Noise. We call the distance between the means of the goal and progress distributions on variable $s$ the Signal to the user’s brain. It describes how far, and in which direction, the user needs to travel along the scale of task outcomes to reach the goal distribution, on average. The stronger the signal, the more stimulus the task provides. However, this also means the more work has yet to be performed by the brain. The reciprocal of this signal represents the $Accuracy$ of the task. The standard deviation of the goal distribution is the tolerance to noise and describes the variation allowed in the goal distribution of the task. We therefore call this the Noise in the task, because it describes how much noise is allowed in the movement towards the goal for distributions to still match. The larger the noise (tolerance), the less precise the user needs to be in executing the task. As such, the reciprocal of this noise represents the $Precision$ of the task. If we return to the example of organizing your socks, the larger the average distance of socks, the Signal, to the drawer, the harder it is to complete the task. The larger the drawer, or Noise tolerance, the easier it will be to complete the task. Fitts’ Law naturally follows from this, but note we are working in bits rather than time. From this also follows that the relative entropy or difficulty of a task is a function of the ratio of the mean distance S to and standard deviation N of the goal. Note that this SNR is different from the one introduced by Shannon as describes a relative signal to a goal.
@@
    KL Divergence. Relative entropy equals the reduction in Bayesian Surprise, in bits, when the user is trying to match progress and goal distributions. This relative entropy is calculated as a Kullback-Leibler Divergence between the goal distribution G and progress distribution P, in bits. Mathematically, this is performed by first converting to bits, subtracting the two distributions (a fraction inside a log is a subtraction), then calculating the surface area using an integral. To describe error, we include an empirical variable $b$ in this equation that relates the KL divergence (where $b=1$) to an observed empirical error rate (see Postulate 10):
@@
    Equal Variance Condition. The KL Divergence can be greatly simplified in cases where the variance of the goal and progress distributions are identical. Bayesian surprise and thus difficulty can be calculated using a simple version of the KL divergence where we take the square of the signal-to-noise ratio, where the signal is the distance between the means of the goal and progress distributions and the noise is the variance in the goal distribution. The mathematical proof of the relative entropy of two Gaussians with equal variance is, in fact, the motivation for the use of squared error in statistics and machine learning (Novak, 2015):
@@
    (14) $=\frac{\left(\mu_{p}-\mu_{g}\right)^{2}+\sigma_{p}^{2}-\sigma_{p}^{2}}{2\ln\left(2\right)\sigma_{g}^{2}}$
@@
    Performance. Human performance, or capacity to process relative entropy, follows the reduction in KL divergence as the user moves along the scale of task outcomes. This capacity to process Bayesian surprise, in bits, follows Shannon-Hartley’s theorem (Equation 2) except with relative entropy. This equals the logarithm of our relative signal-to-noise ratio plus 1.
@@
    Where $b$ is some empirically derived constant describing the average performance. Human performance and error can thus both be described on the same scale of signal-to-noise ratio (SNR). Note that this is a different scale then the task outcome scale. It describes the ratio of the distance between means and standard deviation of the goal. Since performance is a logarithmic function of the SNR, while the amount of Bayesian surprise in a task is a squared function of SNR, users will run out of capacity to process Bayesian surprise as the SNR increases. The user’s Capacity to process Bayesian surprise can be measured, in real time, by calculating the average reduction in KL Divergence per unit of SNR. This capacity equates the amount of cognitive load and/or sensori-motor load induced by a task because it describes the amount of Bayesian surprise processed by the brain. As per Shannon-Hartley’s theorem (Shannon, 1948), optimal performance is defined by the slope or average capacity of the logarithmic function on the SNR scale. Note that we can also express the SNR scale in bits if we take its binary logarithm, in which case capacity becomes a linear function of log SNR.
@@
The ultimate purpose of any user interface design, then, is to achieve Postulate 1 by preventing Postulate 10. Together, our postulates, which are based merely only on derivation and logic, give us a mathematical tool to measure the SNR of a user in real time, i.e., as they perform their task. More importantly, they also a method for measuring the capacity as it relates to potential error, in real time. The easiest way to do this is to measure the mean distance to $\mu_{g}$ over some time window, dividing this by the std. dev, and squaring the result.
Figure 2. Bayesian Fitts’ Law. Distribution of the Posterior probability P(s—o) of the cursor being within the Goal, from top to bottom: in the middle of a trial, at 2/3s of the trial and at the end of a trial. Goal is the 96% area of the Prior probability distribution P(s), on both sides for a reciprocal trial. Note how the velocity of the cursor tracks the uncertainty in cursor position.
[…]
Distribution of the Posterior probability P(s—o) of the cursor being within the Goal, from top to bottom: in the middle of a trial, at 2/3s of the trial and at the end of a trial. Goal is the 96% area of the Prior probability distribution P(s), on both sides for a reciprocal trial. Note how the velocity of the cursor tracks the uncertainty in cursor position.
4. Three Laws of HCI Revisited
@@
4.1. Fitts’ Law
Kording and Wolpert’s work (Körding and Wolpert, 2006; Wolpert, 2007) strongly suggests that sensorimotor control is indeed guided by Bayesian prediction. Wu et al. (Wu et al., 2006) and Williamson et al. (Williamson et al., 2022) also suggest that Fitts’ Law (Fitts, 1954; MacKenzie, 1992) is the result of the sensorimotor system performing Bayesian inference on movement. Figure 2 shows this process. The goal in a Fitts’ Law task is to move the cursor (or hand) reciprocally between two targets or goals at distance A and click within width W. The goal can be modeled as a Prior probability distribution $P(s)$. The 96% area under this probability distribution is sectioned of as a width W that constitutes the visual target. This width W, then, corresponds to a $4\sigma$ variance in the goal. In Fitts’ Law experiments, both the distance A between the means of the priors and their standard deviations, i.e., and width W, are systematically varied (MacKenzie, 1992). The prior distribution is evidenced over successive trials as a histogram of the locations of clicks (Welford, 1968). The figure also shows a cursor. The velocity of this cursor reflects the uncertainty in the position of the cursor relative to the goal $P(s)$. This uncertainty is our progress or posterior distribution $P(s|o)$ and needs to be estimated from observation of the distance of the cursor to the goal given a tolerance. Figure 2 also shows how, when the cursor moves, this velocity and thus the variance in the posterior increases to a maximum uncertainty, after which it decreases to a minimal velocity or uncertainty once the target is reached. On average, however, velocities and thus the variance in the movement must match the variance or size of the priors so as not to click outside the target. In Interactive Inference, movement is modeled by a step-wise iterative update of the posterior to the prior(s), the speed of which is governed by the remaining bits of difficulty during each trial. The calculation of a new estimate of cursor location is modeled as a Bayesian update, in which a new progress distribution is iteratively calculated based on the probability distribution of observations of the hand given the goal distribution (the likelihood, not shown). Indeed, Welford suggested early-on that performance is a function of a signal-to-noise ratio in the brain and that there is a maximum capacity at which the brain can process a Fitts’ Law task due to the “neural noise blurring or distorting signals” (Welford, 1968). MacKenzie (MacKenzie, 1992) suggested the following equation for describing movement time in a Fitts’ Law task:
@@ 4.1. Fitts’ Law
Here, $MT$ equals movemement time, and $a$ and $b$ are empirically determined constants. The equation is modeled after the Shannon-Hartley theorem in Equation 2. However, rather than using the statistical properties of the Bayesian priors, it is based on the physical A and W that form the conditions of the experiment. The binary logarithm component of this equation is known as the Index of Difficulty. It describes, in bits, the difficulty of the task. We propose to rewrite Equation 17 using the statistical properties of Bayesian inference, yielding an equation that describes the Bayesian surprise that is resolved, in bits:
@@
Here, $C$ equals the capacity of the human sensorimotor system in bits, with $a$ and $b$ empirically determined constants. S is the signal in this task, and is given by the mean distance between priors. It is thus equal to A in Equation 17. N describes the noise in this task, and is equal to the standard deviation of the goal distributions. Access to data on the speed of hand movements during Fitts’ Law tasks appears, to our knowledge, limited. This would be needed to investigate the modeling of cursor speed as a posterior update $P(s|o)$. Schmidt et al. (Schmidt et al., 1979) did show the variance in motor impulse in a similar task to be linearly related to the variance in the resulting distribution, i.e., the prior. Their $W_{e}$ metric may well correspond to the $N$ in Equation 18. However, we see further investigation of this relationship as a future direction of this work.
Aside from being more neuromorphic, one advantage of Equation 18 is that capacity can more easily be compared to error, if errors are converted to bits using negative log probability. A second advantage is that it allows real-time measurement of Fitts’ Law task difficulty during a trial by sampling the velocity of the cursor or hand. A third distinct advantage is that it allows the comparison of the capacity of Fitts’ Law tasks with that of other tasks. This is important in cases where the overall capacity of the user’s actions need to be calculated, in bits of unresolved Bayesian surprise. From our earlier discussion follows that once a user runs out of capacity, errors should track a KL divergence after the user runs out of capacity:
@@
Here, $b$ is an empirically determined constant describing the error rate in bits. Because in most experiments the error rate is set to a fixed 4%, beyond which trials are ended (MacKenzie, 1992), there is, at present, limited evidence for such error curve. Welford (Welford, 1968) did demonstrate a method for posthoc correction of goal distributions, so as to adjust for differences in error rates between different devices. His work appears consistent with the above discussion. Wobbrock et al. (Wobbrock et al., 2008) also modeled the probability of an error, with good fit. Our model predicts that user error only tracks the KL divergence once capacity is reached, while their model describes the probability of error rather than the negative log probability. Finally, Guiard et al. (Guiard et al., 2011) analyzed the error in different conditions of a Fitts’ Law experiment. Their RVE metric appear to be consistent with Equation 19.
4.2. Hick-Hyman Law
@@
Where $p_{i}$ is the probability of each choice in the set. Note that as the set gets smaller, the chance of a correct answer increases. As the set gets larger, it decreases by the number of elements in the set. If we examine Equation 7 it is evident that this is equal to a discrete KL divergence for the task, in which the distributions are probability mass functions instead of proper density functions. Here, the probability of the posterior is set to one because there is one item to which the prior needs to be updated. The prior probability is set to $p_{i}$. Conveniently, the binary logarithm of 1 reduces to zero making this relative entropy equal to the total entropy of the task. This means the only relevant parameter in this law is the noise in the prior probabilities: the knowledge the user has at the beginning of the task. According to Wu et al. (Wu et al., 2018), evidence from MRI scans of the brain indeed suggests that the narrowing of the selection from the prior to the observation follows step-wise Bayesian posterior updates. According to them, the cognitive control network in the brain eliminates half the options at each Bayesian update, yielding a binary search with a logarithmic response time. This means Hick-Hyman Law can be rewritten as a capacity function where noise $N$ is a discrete probability that is the reciprocal of the number of choices n:
@@
To our knowledge, there is no experimental evidence of this relationship and we therefore see this as one of the predictions made by our model that requires further empirical investigation.
4.3. Power Law of Practice
@@
Where $a$ and $b$ are empirically derived constants. This equation describes the reduction in entropy, in bits, as an improvement in the signal-to-noise ratio of the task. The time taken per bit then provides the Power Law of Practice. In deep learning, minimizing cross-entropy loss means learning the parameters of a neural network that make the prior distribution as close as possible to the posterior distribution. We propose users likely to do the same when learning a new task.
5. The case of unequal variances
There are some cases in which the simplification of the KL divergence to a square function of the SNR does not work. Consider the example of the distribution of socks that need to be placed in the drawer. We note that we could consider this task a compound task in which we summate the individual KL divergences of the movement of each individual sock. In this example, we are simply going to take the average of these, in an attempt to describe the overall entropy of the task with one KL metric.
@@ 5. The case of unequal variances
5.1. Precision/Speed Trade-off
@@
Variance $\sigma_{g}^{2}$ represents the tolerance in the goal, while variance $\sigma_{p}^{2}$ describes the speed of progress towards that goal, meaning this equation captures the full speed/accuracy trade-off. However, note that the word accuracy is incorrectly used here, and should be worded speed/precision trade-off. The relationship to Fitts’ Law is evident: if the hand moves slower or faster than the size of the target allows, a non-optimal result is produced. Stated more generally, when the user moves, on average, outside the tolerance in the task outcome, the task is executed sub-optimally, and by extension, is performed in non-optimal time or error conditions. In terms of Interactive Inference, the posterior update towards the prior is not of an appropriate size. On average, therefore, $\sigma_{p}^{2}$ should match $\sigma_{g}^{2}$ for optimal results. Hence, the simplified Equation 15 suffices when considering optimal performance as both variances, on average, need to be equal for an optimal speed/precision trade-off. This is the case in Fitts’ Law experiments. Since $\sigma_{p}^{2}$ directly represents velocity when $s$ is a spatial distance metric, this also suggests a linear relationship between velocity and target width (Schmidt et al., 1979).
6. Empirical Study
@@
We now turn to a case study, where we empirically examine whether our general model can predict human performance in a particular situation. If it can, it provides evidence that the general model is worthy of further exploration. In particular, we want to verify, via an empirical study, whether the above postulates could successfully model a task that does not yet have a known model, and function as a real-time measurement tool. We used our simplified KL Divergence to measure the information processing capacity of participants in a simulated target avoidance task: maintaining distance while driving behind a car. Our attempt to develop a better model of a driver’s capacity is more than an academic exercise, as such a model can also impact how modern cars and their interfaces can better recognize and mitigate problems due to driver distraction (Groeger, 2000).
6.1. Study Literature Review

### v4

@@ Introduction
    Difficulty. The difficulty of the task equals the number of bits of relative information entropy remaining to be processed. A fully accomplished task has the minimum bits of relative entropy to be processed (preferably zero). Converting bits to time estimates requires a multiplication of this difficulty with an empirically derived distribution of time performance. E.g., the prior indicates prior statistical knowledge, built up over time, about where socks should be in a second task: that of obtaining a new pair of socks. After the task of organizing socks is complete, when the user wants to obtain a pair of socks, the probability of obtaining them in the prior location is 100%. The posterior distribution of socks corresponds 100% to the prior, the location inside the drawer. By reducing the relative entropy in the distribution of socks in the first task to fit the goal, the user has been able to make the currently estimated location of socks, the posterior or progress distribution P, correspond to the memorized prior, the goal distribution G.
@@
    Goal and Progress Distributions. Goal and progress of a task are defined as probability distributions on the scale of task outcomes $s$. The goal distribution $G$ describes the prior, the probability of positive task outcomes on this scale. The progress distribution $P$ describes the posterior, the probability of being in a current position along the scale of task outcomes $s$. Each distribution has a mean value $\mu$ on this scale. For the goal distribution, the mean $\mu_{g}$ signifies the value with the highest probability of a correct task outcome. For the progress distribution, the mean $\mu_{p}$ denotes the value with the highest probability of the current task outcome. Each distribution also has an associated variance. The variance $\sigma_{g}^{2}$ or standard deviation $\sigma_{g}$ of the goal distribution $G$ describes the tolerance for error that is allowed in achieving goal $G$. The reciprocal of this variance $\frac{1}{\sigma_{g}^{2}}$ is the precision with which the task is to be completed. The variance $\sigma_{p}^{2}$ or standard deviation $\sigma_{p}$ of the progress distribution $P$ describes the speed with which the task outcome is moving towards the goal distribution: the speed of progress on the task. The reciprocal of this variance $\frac{1}{\sigma_{p}^{2}}$ is the precision with which the task is being executed. From postulate 5 it is becomes apparent that, in order for the goal and progress distributions to match, the user must not just reduce the distance between the means of progress and goal distributions to zero, but also match the precision (or variance) of both distributions. That is, a more precise goal will require, at some point during the task, a more precise movement on the scale of possible task outcomes towards that goal. In other words, to avoid error, the user will, at some point, need to move more slowly on the scale of possible task outcomes. This is typically towards the end of the task, when the user is closer to the goal. It also apparent that a not so precise goal definition allows faster progress, i.e., users can be more sloppy in the execution of the task. As such, the ratio between the difference of means and the standard deviation reflects a precision/accuracy trade-off (mistakenly known as speed/accuracy trade-off) that is inherent to all tasks. Because the two distributions will need to match on average, the individual variances of goal and progress distributions can be simplified as equal. This brings us to the next postulate:
@@
    Signal and Noise. We call the distance between the means of the goal and progress distributions on variable $s$ the Signal to the user’s brain. It describes how far, and in which direction, the user needs to travel along the scale of task outcomes to reach the goal distribution, on average. The stronger the signal, the more stimulus the task provides. However, this also means the more work has yet to be performed by the brain. The reciprocal of this signal represents the AccuracyAccuracy of the task. The standard deviation of the goal distribution is the tolerance to noise and describes the variation allowed in the goal distribution of the task. We therefore call this the Noise in the task, because it describes how much noise is allowed in the movement towards the goal for distributions to still match. The larger the noise tolerance, the less precise the user needs to be in executing the task. As such, the reciprocal of this noise represents the PrecisionPrecision of the task. If we return to the example of organizing your socks, the larger the average distance of socks, the Signal, to the drawer, the harder it is to complete the task. The larger the drawer, or Noise tolerance, the easier it will be to complete the task. Fitts’ Law naturally follows from this, but note we are working in bits rather than time. From this also follows that the relative entropy or difficulty of a task is a function of the ratio of the mean distance S to and standard deviation N of the goal. Note that this SNR is different from the one introduced by Shannon as it describes a relative signal to a goal:
@@
    KL Divergence. Relative information entropy measures the reduction in Bayesian Surprise, in bits, when the user is trying to match progress and goal distributions. This relative entropy is calculated as a Kullback-Leibler Divergence between the goal distribution G and progress distribution P, in bits. Mathematically, this is performed by first converting to bits, subtracting the two distributions (a fraction inside a log is a subtraction), then calculating the surface area using an integral. To describe error, we include an empirical variable $b$ in this equation that relates the KL divergence (where $b=1$) to an observed empirical error rate (see Postulate 10):
@@
    Equal Variance Condition. The KL Divergence can be greatly simplified in cases where the goal and progress distributions are normal, with identical variance. Bayesian surprise and thus difficulty can be calculated using a simple version of the KL divergence where we take the square of the signal-to-noise ratio, where the signal is the distance between the means of the goal and progress distributions and the noise is the variance in the goal distribution. The mathematical proof of the relative entropy of two Gaussians with equal variance is, in fact, the motivation for the use of squared error in statistics and machine learning (Novak, 2015):
@@
    (14) $\frac{\left(\mu_{p}-\mu_{g}\right)^{2}+\sigma_{p}^{2}-\sigma_{p}^{2}}{2\ln\left(2\right)\sigma_{g}^{2}}$
@@
    We note that the metric $\frac{\mu_{p}-\mu_{g}}{\sigma_{g}}$ is mathematically identical to Cohen’s d (Cohen, 1988), a well-established statistical measure of effect size between two distributions. Thus, our framework essentially connects this measure to modeling of task difficulty in HCI contexts.
@@
    Performance. Human performance, or capacity to process relative entropy, follows the reduction in KL divergence as the user moves along the scale of task outcomes. This capacity to process Bayesian surprise, in bits, follows Shannon-Hartley’s theorem (Equation 2) but with relative entropy (KL divergence) between prior and posterior, rather than absolute entropy of signals. This equals the logarithm of our relative signal-to-noise ratio plus 1.
@@
    Where $b$ is some empirically derived constant describing the average performance. Human performance and error can thus both be described on the same scale of signal-to-noise ratio (SNR). Note that this is a different scale then the task outcome scale. It describes the ratio of the distance between means and standard deviation of the goal. Since performance is a logarithmic function of the SNR, while the amount of Bayesian surprise in a task is a squared function of SNR, users will run out of capacity to process Bayesian surprise as the SNR increases. The user’s capacity to process Bayesian surprise can be measured, in real time, by calculating the average reduction in KL Divergence per unit of SNR. This capacity equates the amount of cognitive load and/or sensorimotor load induced by a task because it describes the amount of Bayesian surprise processed by the brain. As per Shannon-Hartley’s theorem (Shannon, 1948), optimal performance is defined by the slope or average capacity of the logarithmic function on the SNR scale. Note that we can also express the SNR scale in bits if we take its binary logarithm, in which case capacity becomes a linear function of log SNR.
@@
The ultimate purpose of any user interface design, then, is to allow the user to achieve Postulate 1 while avoiding Postulate 10. Together, our postulates give us a mathematical tool to measure the capacity of the user as it relates to potential error, in real time. The easiest way to do this is to measure the mean distance to $\mu_{g}$ over some time window, dividing this by the std. dev., and squaring the result.
Figure 2. Bayesian Fitts’ Law. Distribution of the Posterior probability P(s—o) of the cursor being within the Goal, from top to bottom: in the middle of a trial, at 2/3s of the trial and at the end of a trial. Goal is the 96% area of the Prior probability distribution P(s), on both sides for a reciprocal trial. Note how the velocity of the cursor tracks the uncertainty in cursor position.:
Three Laws of HCI Revisited
@@
Fitts’ Law
Kording and Wolpert’s work (Körding and Wolpert, 2006; Wolpert, 2007) strongly suggests that sensorimotor control is indeed guided by Bayesian prediction. Wu et al. (Wu et al., 2006) and Williamson et al. (Williamson et al., 2022) also suggest that Fitts’ Law (Fitts, 1954; MacKenzie, 1992) is the result of the sensorimotor system performing Bayesian inference on movement. Figure 2 shows this process. The goal in a Fitts’ Law task is to move the cursor (or hand) reciprocally between two targets or goals at distance A and click within width W. The goal can be modeled as a Prior probability distribution $P(s)$. The 96% area under this probability distribution is sectioned off as a width W that constitutes the visual target. This width W, then, corresponds to a $4\sigma$ variance in the goal. In Fitts’ Law experiments, both the distance A between the means of the priors and their standard deviations, i.e., and width W, are systematically varied (MacKenzie, 1992). The prior distribution is evidenced over successive trials as a histogram of the locations of clicks (Welford, 1968). The figure also shows a cursor. The velocity of this cursor reflects the uncertainty in the position of the cursor relative to the goal $P(s)$. This uncertainty is our progress or posterior distribution $P(s|o)$ and needs to be estimated from observation of the distance of the cursor to the goal given a tolerance. Figure 2 also shows how, when the cursor moves, this velocity and thus the variance in the posterior increases to a maximum uncertainty, after which it decreases to a minimal velocity or uncertainty once the target is reached. On average, however, velocities and thus the variance in the movement must match the variance or size of the priors so as not to click outside the target. In Interactive Inference, movement is modeled by a step-wise iterative update of the posterior to the prior(s), the size of which is governed by the remaining bits of difficulty during the movement. The calculation of a new estimate of cursor location is modeled as a Bayesian update, in which a new progress distribution is iteratively calculated based on the probability distribution of observations of the hand given the goal distribution. Indeed, Welford suggested early-on that performance is a function of a signal-to-noise ratio in the brain and that there is a maximum capacity at which the brain can process a Fitts’ Law task due to the “neural noise blurring or distorting signals” (Welford, 1968). MacKenzie (MacKenzie, 1992) suggested the following equation for describing movement time in a Fitts’ Law task:
@@
Here, MTMT equals movement time, and $a$ and $b$ are empirically determined constants. The equation is modeled after the Shannon-Hartley theorem in Equation 2. We can instead define a KL divergence in the form of Equation 15 in which error $E$ is minimized:
(18) $E=b\,\left(\frac{e+N}{2N}\right)^{2}$
Where NN is the size of the noise tolerance in the cursor (or hand) movement ($4\sigma$ of the goal distribution) and $e$ the distance from current position to the nearest edge of the goal distribution (see Figure 2). If we rewrite this as a gradient descent (i.e., derivative of Equation 18), we obtain an affine update rule that halves the distance of the progress distribution at each step, with identical time intervals $t$:
(19) $e_{t+1}=\tfrac{1}{2}\,e_{t}-\tfrac{1}{2}\,N$
The stop condition is when the cursor finds the near edge $e_{t}\leq 0$. This corresponds to a binary search for the target in which velocity of the cursor is controlled by halving the distance to the far edge of the target, at each time interval, until the near edge is found. During a reciprocal tapping task between the centers of targets, it produces the logarithmic time result found in Equation 17 (Knuth, 1998). In Interactive Inference we formulate this result, instead, in terms of bits updated:
@@
Here, $C$ equals the capacity of the human sensorimotor system (a loop that includes pointing device and cursor) in bits, with $a$ and $b$ empirically determined constants in bits. $S$ is the signal in this task, and is given by the distance between the center of the goals. Because Equation 19 describes the movement control function, rather than the exact end of movement, over many reciprocal tapping trials the expected end location clusters around the center–to–center distance $S$, with a slight bias towards the near edges. NN describes the noise tolerance of the cursor (or hand) in this task, and is equal to the width of the goal distributions. The main benefit of Equation 19 is that it can be used to evaluate Bayesian surprise, in bits, during the pointing task. Empirical validation of this Bayesian underpinning of Fitts’ Law remains ongoing. Schmidt et al. (Schmidt et al., 1979) did show the variance in motor impulse in a similar task to be linearly related to the variance in the resulting distribution, i.e., the prior. Their $W_{e}$ metric may correspond to the NN in Equation 20.
Another advantage of the above formulation is that capacity can more easily be compared to error, if errors are converted to bits using negative log probability. A third advantage is that it may allow the comparison of the capacity of Fitts’ Law tasks with that of other tasks. This is important in cases where the overall capacity of the user’s actions needs to be calculated, in bits of unresolved Bayesian surprise. From our earlier discussion follows that once a user runs out of capacity, errors should track a KL divergence after the user runs out of capacity:
@@
Here, $b$ is an empirically determined constant describing the error rate in bits. There is some evidence for such error curve (de la Fuente, 2014). Welford (Welford, 1968) did demonstrate a method for posthoc correction of goal distributions, so as to adjust for differences in error rates between different devices. His work appears consistent with the above discussion. Wobbrock et al. (Wobbrock et al., 2008) also modeled the probability of an error, with good fit. Our model predicts that user error only tracks the KL divergence once capacity is reached, while their model describes the probability of error rather than the negative log probability. Finally, Guiard et al. (Guiard et al., 2011) analyzed the error in different conditions of a Fitts’ Law experiment. Their RVE metric also appears to be consistent with Equation 21.
Hick-Hyman Law
@@
Where $p_{i}$ is the probability of each choice in the set. Note that as the set gets smaller, the chance of a correct answer increases. As the set gets larger, it decreases by the number of elements in the set. If we examine Equation 7 it is evident that this is equal to a discrete KL divergence for the task, in which the distributions are probability mass functions instead of proper density functions. Here, the probability of the posterior is set to one because there is one item to which the prior needs to be updated. The prior probability is set to $p_{i}$. Conveniently, the binary logarithm of 1 reduces to zero making this relative entropy equal to the total entropy of the task. This means the only relevant parameter in this law is the noise in the prior probabilities: the knowledge the user has at the beginning of the task. According to Wu et al. (Wu et al., 2018), evidence from MRI scans of the brain indeed suggests that the narrowing of the selection from the prior to the observation follows step-wise Bayesian posterior updates. According to them, the cognitive control network in the brain eliminates half the options at each Bayesian update, yielding a binary search with a logarithmic response time. This means Hick-Hyman Law can be rewritten as a capacity function where noise NN is a discrete probability that is the reciprocal of the number of choices n:
@@
We see this as one of the predictions made by our model that requires further empirical investigation.
Power Law of Practice
@@
Where $a$ and $b$ are empirically derived constants. This equation describes the reduction in entropy, in bits, as an improvement in the signal-to-noise ratio of the task. The time taken per bit then provides the Power Law of Practice.
The Case of Unequal Variances
There are some cases in which the simplification of the KL divergence to a square function of the SNR does not work. Consider the example of the distribution of socks that need to be placed in the drawer. We note that we could consider this task a compound task in which we sum the individual KL divergences of the movement of each individual sock. In this example, we are simply going to take the average of these, in an attempt to describe the overall relative entropy of the task with one KL metric.
@@
Precision/Speed Trade-off
@@
Variance $\sigma_{g}^{2}$ represents the tolerance in the goal, while variance $\sigma_{p}^{2}$ describes the speed of progress towards that goal, meaning this equation captures the full speed/accuracy trade-off. However, note that the word accuracy is incorrectly used here, and should be worded speed/precision trade-off. The relationship to Fitts’ Law is evident: if the hand moves slower or faster than the size of the target allows, a non-optimal result is produced. Stated more generally, when the user moves outside the tolerance in the task outcome, the task is executed sub-optimally, and by extension, is performed in non-optimal time or error conditions. In terms of Interactive Inference, the posterior update towards the prior is not of an appropriate size. On average, therefore, $\sigma_{p}^{2}$ should match $\sigma_{g}^{2}$ once the goal is approached. Hence, the simplified Equation 15 suffices when considering optimal performance as both variances, on average, need to be equal for an optimal speed/precision trade-off. This appears to be the case in Fitts’ Law. Since $\sigma_{p}^{2}$ directly represents velocity when $s$ is a spatial distance metric, this also suggests a linear relationship between velocity and target width (Schmidt et al., 1979).
Empirical Study
@@
We now turn to a case study, where we empirically examine whether our general model can predict human performance in a particular situation. If it can, it provides evidence that the general model is worthy of further exploration. In particular, we want to verify, via an empirical study, whether the above postulates could successfully model a task that does not yet have a known model, and function as a real-time measurement tool. We used our simplified KL Divergence to measure the information processing capacity of participants in a simulated target avoidance task: maintaining distance while driving behind a car. Our attempt to develop a better model of a driver’s capacity is more than an academic exercise, as such a model can help modern car interfaces better recognize and mitigate problems of driver distraction (Groeger, 2000).
Study Literature Review